an investor purchased 100 shares of fifth third bank stock and 100 shares of santee electric cooperative stock.the probability the bank stock will appreciate over a year is 70. the probability the electric utility will increase over the same period is 60. what is the probability both stocks appreciate during the period?
what is the probability the bank appreciates but the utility does not?
what is the probability at least one of the stocks appreciates?

Answer by ikleyn(52788)

.

(a)   What is the probability both stocks appreciate during the period?
     P = 0.7*0.6.


(b)   What is the probability the bank appreciates but the utility does not?
     P = 0.7*(1-0.6).


(c)   what is the probability at least one of the stocks appreciates?
     P = 0.7*(1-0.6) + (1-0.7)*0.6


Solved, answered and completed.